Find the volume generated by rotating the region bounded by y = x, x = 1 and y² = 4x, about the x axis.
1
Expert's answer
2020-04-06T13:02:15-0400

Consider a region that is bounded by two curves "y=f(x)" and "y=g(x)" , between "x=a" and "x=b."

("f(x), g(x)" are continuous and non-negative on the interval "[a,b]" and "f(x)\leq g(x)" )

The volume of the solid formed by revolving the region about the x-axis is

"V=\pi \int \limits_a^b\big( [f(x)]^2 -[g(x)]^2\big) dx"


We have three curves: "y=x, y^2=4x" and "x=1" .

The region that is bounded by them can be defined as a region that is bounded by "f(x)=2\sqrt {x}" and "g(x)=x" , between "x=0" and "x=1" .

So, the volume generated by rotating this region is:

"V=\pi \int \limits_0^1\big( (2\sqrt{x})^2 -x^2\big) dx= \pi \int \limits_0^1\big( 4x-x^2\big) dx=
\pi \big(2x^2-\frac{x^3}{3}\big) \big|_0^1=\pi (2-\frac{1}{3})=\frac{5\pi}{3}"


Answer: the volume is "\frac{5\pi}{3}."